Thank you for helping me! The final question on my homework task is exceedingly hard and I have been trying for hours to get the answer.
Here it is: Adam has sticks of the following lengths: (3x+4)cm (2x+10)cm (x+12)cm
He puts all 3 sticks together to make a triangle. The triangle is Isosceles. Calculate the three possible values of X.
Please reply as soon as possible, as this homework is due tomorrow.
Thanks

Answer by mananth(16946)

(3x+4)cm (2x+10)cm (x+12)
option I
(3x+4)cm =(2x+10)cm
3x-2x=6
x=6
Option II
(3x+4)= (x+12)
3x-x=12-4
2x=8
x=4
Option III
(2x+10)= (x+12)
2x-x=12-10
x=2